# TranDRL: A Transformer-Driven Deep Reinforcement Learning Enabled Prescriptive Maintenance Framework

### v1

@@ IV Proposed Framework
1 Initialize the positional encoding $\mathbf{PE}$;
2 Initialize the input embeddings $\mathbf{E}$;
@@
4       Apply multi-head attention with residual connections and layer normalization;
5       Apply position-wise feedforward networks with residual connections and layer normalization;
@@
8Implement a linear transformation and a softmax activation function;
Input: Observation $s$, exploration rate $\epsilon$, learning rate $\alpha$, discount factor $\gamma$
Output: Action $a$
9 Initialize the neural network weights;
10 for each episode do
11       Initialize the starting state $s$;
12       while $s$ is not a terminal state do
13             Select action $a$ using an $\epsilon$-greedy policy;
14             Execute action $a$, observe reward $r$ and successor state $s^{\prime}$;
15             Update neural network weights utilizing the Temporal Difference (TD) error $\delta$ and the learning rate $\alpha$;
16             Adjust the exploration rate $\epsilon$;
17             Update current state $s\leftarrow s^{\prime}$;
18
19       end while
20
@@ 17             Update current state $s\leftarrow s^{\prime}$;
Algorithm 1 The Transformer and Deep Reinforcement Learning Enabled Prescriptive Maintenance Framework
IV-A Remaining Useful Life Prediction Module
Transformers, rooted in Natural Language Processing with models like BERT, GPT-2, and T5 [17], have become valuable tools in time-series data analysis. This is largely due to their proficiency in understanding complex temporal sequences. Our model harnesses the Transformer’s capabilities to forecast the Remaining Useful Life (RUL) from historical time-series sensor data. At the heart of this approach lies the self-attention mechanism, which prioritizes different time steps based on their importance, adapting finely to the intricacies of the data.
IV-A1 Encoder Module
For a given time-series $x=\{x_{1},x_{2},...,x_{T}\}$, the Encoder Module converts it into hidden states $h=\{h_{1},h_{2},...,h_{T}\}$. Each Transformer block in this module comprises:
  • Multi-Head Self-Attention Mechanism: Given matrices for query $Q$, key $K$, and value $V$, the attention mechanism is:
    $\text{Attention}(Q,K,V)=\text{Softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V,$
    with $d_{k}$ being the dimension of queries and keys.
  • Feed-Forward Neural Network: To capture non-linear patterns, this is represented as:
    $FFN(x)=\max(0,xW_{1}+b_{1})W_{2}+b_{2},$ (1)
    with weight matrices $W_{1},W_{2}$ and bias terms $b_{1},b_{2}$.
IV-A2 Self-Attention Mechanism
This module facilitates the model in emphasizing or downplaying certain temporal characteristics. The attention weights $\alpha$ compute the weighted sum of features:
$\alpha_{ij}=\text{Softmax}\left(\frac{Q_{i}K_{j}^{T}}{\sqrt{d_{k}}}\right),$
IV-A3 Decoder Module
Post the encoding of time-series data into hidden states, the decoder refines this information to produce outputs conducive for RUL prediction. Given hidden states $h$, the Decoder Module yields output states $o=\{o_{1},o_{2},...,o_{T}\}$. Each block in the decoder encompasses:
  • Multi-Head Cross-Attention Mechanism: Attending to the encoder’s outputs while forming the decoder’s states, it’s represented as:
    $\text{Attention}(Q_{d},K_{e},V_{e})=\text{Softmax}\left(\frac{Q_{d}K_{e}^{T}}{\sqrt{d_{k}}}\right)V_{e},$
    with $d_{k}$ being the dimension of the queries and keys.
  • Feed-Forward Neural Network: To discern patterns in the decoder states:
    $FFN(o)=\max(0,oW_{3}+b_{3})W_{4}+b_{4},$
    with weight matrices $W_{3},W_{4}$ and biases $b_{3},b_{4}$.
IV-A4 Prediction Module
At the culmination of our model, the prediction module utilizes the hidden states $h$ to estimate the RUL. This involves one or several dense layers, expressed as:
$\text{RUL}=\sigma(hW+b),$
with weight matrix $W$, bias $b$, and activation function $\sigma$ ensuring the RUL remains within the intended range.
IV-B Maintenance Action Recommendation Module
Upon retrieving the RUL predictions from the previous module, the maintenance action recommendation module initiates. This module uses the predicted RUL as a primary determinant, facilitating the calibration of the urgency, type, and extent of necessary maintenance interventions. However, an extended predicted RUL diminishes the immediate demand for aggressive maintenance interventions. In contrast, a reduced RUL necessitates prompt preventive or remedial actions. Recommendations from this module undergo further refinement, adapting to specific industrial settings and the complexities inherent to the particular equipment or system.
This module predominantly employs two methodologies:
@@ IV-B Maintenance Action Recommendation Module
  • Rule-Based Systems: These systems are grounded in predetermined heuristic principles. As an instance, should the RUL prediction for a component fall below a designated limit (e.g., 10 hours), an immediate maintenance notification is dispatched. On the other hand, if the RUL surpasses a set benchmark, periodic inspections may be deemed adequate.
@@
  • Machine Learning Classifiers: These classifiers are cultivated using historical datasets, where earlier RUL predictions are coupled with their corresponding effective maintenance interventions. Inputting the present RUL prediction enables the classifier to recommend the most probable efficacious maintenance strategy, grounded in historical insights.
The harmonious integration of the RUL predictions with the Maintenance Action Recommendation Module ensures that maintenance schedules are both anticipatory and informed by data, culminating in augmented system dependability, extended equipment longevity, and streamlined operational expenditures.
Upon acquiring the RUL predictions from the antecedent module, the Maintenance Action Recommendation Module becomes operational. The forecasted RUL, indicative of a system’s operational state, is assimilated into the state representation within the Markov Decision Process (MDP) framework. Subsequently, this representation guides the meticulous formulation of optimal maintenance strategies over time.
IV-B1 Markov Decision Process (MDP) for Maintenance
@@
  • Environment State Space. As established, the system can be in one of $n$ distinct degradation states, with $\mathbf{S_{t}}=[S_{1}^{t},S_{2}^{t},\ldots,S_{n}^{t}]$ encapsulating all possible states the system can take at time $t$.
@@
  • Action Space. If we denote the number of possible maintenance actions by $m$, then the action space $A$ can be represented as $A=\{a_{1},a_{2},\ldots,a_{m}\}$.
  • State Transition Probabilities. The probability of transitioning from state $s_{i}^{t}$ to state $s_{j}^{t+1}$ given action $a_{k}^{t}$ is represented by:
    $P_{i,j}^{k}=P(s_{t+1}=s_{j}^{t+1}|s_{t}=s_{i}^{t},a_{t}=a_{k}^{t}).$
@@
    The complete state transition probability matrix given action $a_{k}^{t}$ would then be a matrix $\mathbf{P^{k}}$ where each entry $P_{i,j}^{k}$ represents the transition probability from state $s_{i}^{t}$ to state $s_{j}^{t+1}$.
@@
    where $\gamma$ is the discount factor.
  • Bellman Optimality Equation. The value of a state under an optimal policy $\pi^{*}$ obeys the recursive relationship:
@@
V Experiments
In this section, we evaluate our proposed framework on the real-world dataset. Experiments are done on the XXX.
@@ V-A Dataset
The NASA’s C-MAPSS dataset [18] is partitioned into four subsets denoted as FD001, FD002, FD003, and FD004, each having a training and test set, reflecting distinct operating and fault modes. The training data comprises inputs from three operating conditions and 21 sensors’ measurements. For training, every engine data point serves as a sample paired with its RUL. We applied a piecewise linear degradation model [19] for RUL determinations. In testing, the last cycle data point for each engine generally stands as the sample, with actual RUL values included in the dataset.
V-B Evaluation Metric
To assess the model’s predictive accuracy, we utilized the Mean Squared Error (MSE), given by
$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2},$ (2)
which quantifies the average squared differences between predicted $\hat{y}_{i}$ and observed $y_{i}$ values. Additionally, the coefficient of determination, $R^{2}$, defined as
$R^{2}=1-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}},$ (3)
where $\bar{y}$ is the mean of observed values, was employed to measure the proportion of variance explained by the model. A higher $R^{2}$ suggests a better fit to the observed data.
V-C Remaining Useful Life Prediction
@@
The Temporal Fusion Transformer (TFT) [20] addresses the limitations of traditional time series forecasting models by offering enhanced interpretability and the capability to deal with multiple prediction horizons. It merges the advantages of Recurrent Neural Networks (RNNs) and Transformers [21] into a unified encoder-decoder architecture.
@@
Encoder: The encoder transforms historical time-series data into a high-dimensional space using a blend of convolutional layers and self-attention mechanisms, thereby learning rich representations of the input.
@@
Temporal Attention: This specialized module empowers the model to focus on different time-steps variably, facilitating the capture of dynamic temporal patterns and long-range dependencies.
@@
Decoder: Tailored for multi-horizon prediction, the decoder employs an autoregressive model that recursively uses its previous forecasts as inputs for subsequent steps.
@@
Static Inputs: Apart from time-dependent features, TFT can also integrate static inputs to provide a comprehensive context for forecasting.
@@
Interpretable Components: To ensure transparency, TFT integrates mechanisms like gating and linear autoregressive decoders that allow for human-readable prediction interpretation.
Figure 1: FD001.
Figure 2: FD002.
Figure 3: FD003.
Figure 4: FD004.
V-D Optimal Maintenance Actions Recommendation
@@
In the previous stage, we succeed in predicting RUL using four datasets from NASA C-MAPSS. But results from the prediction cannot provide enough information for engineers to make decisions on which maintenance actions to take. Thus, we may recommend corresponding actions based on the cost and benefits. To optimize results, we leverage DRL algorithms, which helps to find the best trade-off.
V-D1 States and Actions
@@
The state space is generated from PCA features of the RUL resulted from the prediction outcomes of the NASA’s C-MAPSS Aircraft Engine Simulator dataset. These PCA features provide a compact yet informative snapshot of the machine’s current health condition. The action space contains three primary actions:
@@
  • Partial Maintenance ($a_{2}$), suitable for moderate wear states, involves preventive actions like lubrication to extend machine life at moderate costs;
  • Complete Overhaul ($a_{3}$), necessary for critical machine states, entails comprehensive repairs and replacements to avert catastrophic failures but incurs high costs and downtime.
@@
V-D2 Probabilities Model
In our study, the transition and action cost matrices are empirically calibrated using the NASA C-MAPSS dataset. Specifically, each entry in the state transition matrix $p(S_{t+1}|S_{t},a_{t})$ reflects the probability of transitioning from state $S_{t}$ to $S_{t+1}$ when action $a_{t}$ is taken, and these probabilities are derived through statistical analyses of state sequences and their corresponding RUL values in the C-MAPSS dataset. Moreover, action-related costs, such as maintenance expenses or energy consumption, are grounded in real-world data from this dataset. In this MDP framework, the transition matrix has dimensions $(10,3,10)$, representing 10 states, 3 actions, and 10 possible next states, all closely mirroring scenarios that commercial aero-engines might encounter.
V-D3 Reward Model
In predictive maintenance applications using DRL, the objective is not only to maximize the RUL of machinery but also to minimize the operational costs and downtime. To fulfill these multi-objective requirements, we propose a reward function that combines these elements in a weighted manner. The reward function $R(S_{t},a_{t},S_{t+1})$ is formulated as:

### v2

@@ IV-A4 Prediction Module
1 Initialize the positional encoding;
2 Initialize the input embeddings;
@@
       /* Encoder Layer $i$/
4       Apply multi-head self-attention to the input sequence;
5       Add residual connection followed by layer normalization;
6       Apply position-wise feed-forward networks;
7       Add another residual connection followed by layer normalization;
8
9 end for
/* Decoder */
10 if task requires a decoder then
11       for $j\leftarrow 1$ to $L$ do
             /* Decoder Layer $j$/
12             Apply masked multi-head self-attention to the output of the encoder;
13             Add residual connection followed by layer normalization;
14             Apply multi-head attention combining encoder outputs and decoder inputs;
15             Add residual connection followed by layer normalization;
16             Apply position-wise feed-forward networks;
17             Add another residual connection followed by layer normalization;
18
19       end for
20
21 end if
22Generate output sequence $\mathbf{y}$ based on the final layer’s output;
return $\mathbf{y}$
Algorithm 2 A Transformer algorithm for a single machine.
IV-B Case 2: A Federated RUL Prediction Module for Multi-Machines
The multi-machine case is featured by a federated model, which dispatches an individual training model for similar machines. The local model takes as input the local time series data $\mathbf{X}$ and outputs an estimation of the RUL $R(t)$. By combining the strengths of multiple models extracted from different machines, the approach improves the reliability and accuracy of RUL predictions, enabling industries to make informed decisions regarding maintenance schedules and machinery operations. The averaged model for RUL prediction leverages an ensemble of predictive models to provide a more precise estimate of RUL, enhancing the efficiency and reliability of predictive maintenance in industries.
As shown in the Algorithm 3, given a set of time series data $\mathbf{X}$, the ensemble model aims to predict the RUL $R(t)$ of multiple machines. Formally, the ensemble model can be defined as a function:
$R(t)=f_{\text{ensemble}}(\mathbf{X}),$ (4)
where $f_{\text{ensemble}}$ is the ensemble model that aggregates the predictions of multiple individual models.
By integrating the predictions of multiple models, the ensemble approach refines the reliability and accuracy of the RUL predictions. This enhanced prediction quality allows industries to make precise maintenance schedules and optimize machine operations. The averaged model for RUL prediction can be represented as:
$R(t)=\frac{1}{m}\sum_{i=1}^{m}f_{i}(\mathbf{X}),$ (5)
where $m$ is the number of models in the ensemble, and $f_{i}$ represents the individual predictive model. This averaging method capitalizes on the collective strength of diverse predictive algorithms. Based on the predictive RULs, the infrastructure can conduct equipment maintenance or replacement to ensure uninterrupted and reliable functionality of the machines.
Input: Distributed input sequences $\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{K}$ from $K$ machines
Output: Aggregated output sequences $\mathbf{y}_{1},\mathbf{y}_{2},\ldots,\mathbf{y}_{K}$
1 Initialize a central model with positional encoding and input embeddings; for each training round do
       /* Local Training on Each Machine */
2       for each machine $k=1$ to $K$ do
3             Receive model parameters from the central model;
4             Apply Algorithm 2 with the given number of layers on local sequence $\mathbf{x}_{k}$;
5             Compute local updates and send them back to the central model;
@@
      /* Aggregation at the Central Model */
8       Aggregate local updates from all machines;
9       Update central model parameters;
10
@@ 9       Update central model parameters;
/* Output Generation */
12 for each machine $k=1$ to $K$ do
13       Send final model parameters to machine $k$;
14       Generate output sequence $\mathbf{y}_{k}$ using the final model on local input $\mathbf{x}_{k}$;
15
16 end for
return $\mathbf{y}_{1},\mathbf{y}_{2},\ldots,\mathbf{y}_{K}$
Algorithm 3 A Federated Transformer Algorithm for Multiple Machines
The distinction between Case 1 and Case 2 lies primarily in their scale and data handling methods. Case 1 focuses on predicting the RUL for a single machine using a Transformer model, which excels in interpreting complex temporal sequences in sensor data. This approach is highly specialized for analyzing the intricate patterns of an individual machine. In contrast, Case 2 extends to multiple machines, employing a federated learning approach with an ensemble model. This method aggregates data from various machines, synthesizing diverse inputs for a broader analysis. It suits for a industry which contains multiple similar production lines or machines. Additionally, Case 2 could apply Case 1’s algorithm to each machine and then focus on multi-machine data aggregation.
IV-C Maintenance Action Recommendation Module
Upon obtaining the RUL predictions from the previous module, the maintenance action recommendation module initiates. This module uses the predicted RUL as a primary determinant, facilitating the calibration of the urgency, type, and extent of necessary maintenance interventions. If the predicted RUL of a machine or component is longer, there is less need for urgent maintenance. In contrast, if the predicted RUL is shorter, it is important to quickly perform maintenance actions or repairs to prevent problems. Thus, we select the most important part of RUL using principal component analysis (PCA) as follows:
@@ IV-C Maintenance Action Recommendation Module
Rule-Based Systems. These systems are established on the basis of predetermined heuristic principles. For instance, should the RUL prediction for a component fall below a designed limit (e.g., 10 hours), an immediate maintenance notification is dispatched. On the other hand, if the RUL surpasses the limit, periodic inspections is considered sufficient.
@@
Strategy-Learning Agents. These agents are trained using historical datasets, where earlier RUL predictions are coupled with their corresponding effective maintenance interventions. Inputting the present RUL prediction enables the agent to recommend the most probable efficient maintenance actions, grounded in historical insights.
The appropriate integration of the RUL predictions with the maintenance action recommendation module ensures that maintenance schedules is efficient and data-driven, ultimately leading to improved system reliability, longer equipment lifespan, and reduced operational costs. Thus, the predicted RUL, indicative of a system’s operational state, is assimilated into the state representation within a Markov Decision Process (MDP)-based framework. The subsequent discussions elaborate the maintenance strategy-learning procedure under the MDP framework, which includes following main components:
@@
Environment State Space. As established, the system can be in one of $n$ distinct degradation states, with $\mathbf{S}^{t}=[S_{1}^{t},S_{2}^{t},\ldots,S_{n}^{t}]$ consisting all possible $n$ states that the system can take at time $t$.
@@
Action Space. If we denote the number of possible maintenance actions by $m$, the action space $\mathcal{A}$ can be represented as $\mathcal{A}=\{a_{1},a_{2},\ldots,a_{m}\}$.
@@
State Transition Probabilities. The probability of transitioning from state $s_{i}^{t}$ to state $s_{j}^{t+1}$ given action $a_{k}^{t}$ is represented by $P_{i,j}^{k}=P(s_{t+1}=s_{j}^{t+1}|s_{t}=s_{i}^{t},a_{t}=a_{k}^{t}),$ where the complete state transition probability matrix given action $a_{k}^{t}$ would then be a matrix $\mathbf{P}^{k}$ where each entry $P_{i,j}^{k}$ represents the transition probability from state $s_{i}^{t}$ to state $s_{j}^{t+1}$.
@@
where $\gamma$ is the discount factor for future rewards.
Optimal Policies under Bellman Equation. The value of a state under an optimal policy $\pi^{*}$ obeys the following recursive Bellman optimality equation
@@
On top of this MDP-based policy derivation model, we are able to develop a DRL-based plug-and-play module for maintenance action recommendation using neural network-based function/policy approximation, by accommodating a number of state-of-the-art DRL algorithms. Algorithm 4 shows the details of leveraging different DRL algorithms, including DQN, SAC, PPO, for selecting optimal maintenance action $a$ through evaluating the performance of each policy. Details of hyperparameters setting and how to apply the algorithm to the real-world case are explained in Section V-C4.
Input: Exploration rate $\epsilon$ (for DQN, SAC), learning rate $\alpha$, discount factor $\gamma$ and policy update frequency (for PPO, SAC)
Output: Optimal action $a$ for maintenance
/* Initialization */
1 Initialize the policy network and optionally value/critic network(s) with the input hyperparameters;
2 Initialize target network(s) if applicable (e.g., for DQN, SAC);
3 Initialize replay buffer if applicable;
4 for each episode do
5       Initialize the starting state $\bm{s}$;
6       while $\bm{s}$ is not a terminal state do
             /* Action Selection */
7             if using DQN or SAC then
8                   Select action $a$ using an $\epsilon$-greedy policy or according to the current policy and exploration strategy;
9
10             else if using PPO then
11                   Select action $a$ according to the current policy;
12
13             end if
14            Execute action $a$, observe reward $r$ and successor state $\bm{s^{\prime}}$;
             /* Policy and (or) Value Network Update */
15             if using DQN then
16                   Store transition $(\bm{s},a,r,\bm{s^{\prime}})$ in replay buffer;
17                   Sample mini-batch and update the network(s) using TD error;
18                   Periodically update target network;
19
20             else if using PPO then
21                   Accumulate data for policy update;
22                   Update policy using PPO;
23
24             else if using SAC then
25                   Store transition in replay buffer;
26                   Sample mini-batch and update policy and value networks using SAC update rules;
27                   Adjust temperature parameter for entropy maximization;
28
29             end if
            /* Exploration Rate Adjustment */
30             Adjust the exploration rate $\epsilon$ if applicable;
31             Update current state $\bm{s}\leftarrow\bm{s^{\prime}}$;
32
33       end while
      /* Policy Evaluation */
34       Periodically evaluate the performance of the current policy;
35
36 end for
return Optimal maintenance action $a$
Algorithm 4 A Deep Reinforcement Learning Algorithm for Maintenance Action Recommendation
V Case Study
In this section, we perform a case study on aircraft engine maintenance using NASA’s C-MAPASS dataset [28] to evaluate our proposed TranDRL framework.
@@ V-A Dataset
The NASA’s C-MAPSS dataset [28] contains four subsets denoted as FD001, FD002, FD003, and FD004, each having a training and test set, reflecting distinct operating and fault modes. The training data comprises inputs from three operating conditions and 21 sensors’ measurements. For training, every engine data point serves as a sample paired with its RUL. In the experiment, we applied a piecewise linear degradation model [29] for RUL determinations. The last cycle data point for each engine generally stands as the sample, with actual RUL values included in the dataset.
Figure 2: An Overview of the Temporal Fusion Transformer.
V-B Remaining Useful Life Prediction
@@
The framework given in Algorithm 3 makes it suitable to incorporate the Transformer-based time-series event forecasting networks with other feature extraction and post-processing modules. In the case study, we adopt the Temporal Fusion Transformer (TFT) [30] in Algorithm 3 to utilizes the advantages of both recurrent neural networks and Transformers [31] into a unified encoder-decoder architecture. Compared to traditional time series forecasting models, e.g., Long Short-Term Memory networks (LSTMs) [32], TFT offers enhanced interpretability and the capability to deal with multiple prediction horizons. The structure of TFT is shown in the Figure 2 and the includes following components:
@@
Encoder. The encoder transforms the raw time-series data into a high-dimensional space using convolutional layers and self-attention mechanisms, thereby learning rich representations of the input.
@@
Temporal Attention. This specialized module let the model focus on different time-steps, facilitating the capture of dynamic temporal patterns and long-range dependencies.
@@
Decoder. Tailored for multi-horizon prediction, the decoder employs an autoregressive model that recursively uses its previous forecasts as inputs for forecasting subsequent steps.
@@
Static Inputs. Apart from time-dependent features, TFT can also integrate static inputs to provide a comprehensive context for forecasting.
@@
Interpretable Components. To ensure transparency, TFT integrates mechanisms like gating and linear autoregressive decoders that allow for human-readable prediction interpretation.
Evaluation Results. In our comparisons, Figures 3 to 6 demonstrate the TFT algorithm’s superior performance over NBeats [33] and DeepAR [34] in predicting the RUL for a individual machine, as evidenced by consistently higher RUL values. This suggests TFT’s enhanced capability in capturing complex, time-dependent patterns in sensor data, highlighting its robustness across various machine behaviors. In contrast, Figure 7 shows that while federated learning models generate reasonably good average RUL values across multiple machines, they may not capture the intricate details as effectively as single-machine models. However, Figures 8 to 11 show a significant bias when models trained on a single machine are applied to other machines, underscoring the challenge of model generalizability in diverse operational environments. This analysis illustrates the trade-off between the granularity and specificity of single-machine models and the broader applicability but potentially lower precision of federated learning approaches, highlighting the need for context-aware modeling strategies in predictive maintenance.
Figure 3: The RUL prediction result for machine FD001.
Figure 4: The RUL prediction result for machine FD002.
Figure 5: The RUL prediction result for machine FD003.
Figure 6: The RUL prediction result for machine FD004.
Figure 7: Comparison of RUL prediction results of four machines using the averaged model.
Figure 8: Comparison of RUL prediction results of four machines using the model trained from machine FD001.
Figure 9: Comparison of RUL prediction results of four machines using the model trained from machine FD002.
Figure 10: Comparison of RUL prediction results of four machines using the model trained from machine FD003.
Figure 11: Comparison of RUL prediction results of four machines using the model trained from machine FD004.
V-C Optimal Maintenance Actions Recommendation
@@
In the RUL prediction stage, experiment results show that our proposed framework succeeds in predicting RULs using four datasets from NASA C-MAPSS dataset [28]. However, a gap still exists between the prediction results and the maintenance action decisions needed by engineers. To make action recommendation by weighing the cost and benefits, we leverage DRL algorithms as given in Algorithm 4, which helps to find the best trade-off.
V-C1 States and Actions
@@
The state space is generated from PCA features of the RUL resulted from the prediction outcomes of the NASA’s C-MAPSS Aircraft Engine Simulator dataset. These PCA features provide a compact and important part of the machine’s current health condition. The action space contains three actions:
@@
  • Partial Maintenance ($a_{2}$), suitable for moderate unhealthy states, involves preventive actions like lubrication to extend machine life at moderate costs;
  • Complete Overhaul ($a_{3}$), necessary for critical machine states, entails comprehensive repairs and replacements to prevent catastrophic failures but incurs high costs and downtime.
@@
V-C2 Probabilities Model
In our study, the transition and action cost matrices are empirically calibrated using the NASA C-MAPSS dataset. Specifically, each entry in the state transition matrix $p(S_{t+1}|S_{t},a_{t})$ reflects the probability of transitioning from state $S_{t}$ to $S_{t+1}$ when action $a_{t}$ is taken, and these probabilities are derived through statistical analyses of state sequences and their corresponding RUL values in the C-MAPSS dataset. Moreover, action-related costs, such as maintenance expenses or energy consumption, are grounded in real-world data from this dataset. In this MDP framework, the transition matrix has a dimension $(10,3,10)$, representing transitions between $10$ discretized states given $3$ actions, all closely reflecting scenarios that commercial aero-engines might encounter.
V-C3 Reward Model
The objective of maintenance actions recommend algorithms is not only to maximize the RUL but also to minimize the operational costs and downtime. To fulfill these multi-objective requirements, we propose a reward function that combines these elements in a weighted manner. The reward function $R(S_{t},a_{t},S_{t+1})$ is formulated as: